OpenStudy (anonymous):

NEED HELP!! what number is 7%of 128?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

lalaly what answer you got???

OpenStudy (anonymous):

8.96 <-- i think its wrong

OpenStudy (saifoo.khan):

\[\frac{128}{7} \times 100\]

OpenStudy (anonymous):

.07 * 128 = 8.96

OpenStudy (anonymous):

\[\frac{7}{100}*128\]

OpenStudy (anonymous):

so wait am i right or wrong

OpenStudy (mimi_x3):

Use a calculator its common sense

OpenStudy (anonymous):

The answer is 8.96. If that was your answer, you were right. Otherwise, you were wrong. Have a nice day. :)

OpenStudy (anonymous):

i did and got 8.96

OpenStudy (anonymous):

yes,you are right

OpenStudy (saifoo.khan):

Right.

OpenStudy (saifoo.khan):

Ajmal & xen in right.

OpenStudy (anonymous):

ok thanks for clearing my doubts xD